# A Comprehensive Survey on Knowledge Distillation

### v1

@@ II-C Similarity-based Distillation
One of the pioneering works was [52], which introduced the Flow of Solution Procedure (FSP). This method involved computing the inner product between features from two layers in the teacher network and defining the squared L2 norm as the cost function for each pair of layers between the teacher and student networks. In a related work, [44] suggested distilling the relative dissimilarity between intermediate representations of different examples. Subsequent studies proposed distilling the 8-neighborhood similarity of logits [84] and distilling similarity of features and logits among multiple teachers [85].
@@
Recently, similarity distillation has garnered increased attention, with numerous methods proposed, each focusing on different levels of similarity such as instance/sample-level similarity [45, 25, 46, 86, 47, 50, 51, 49, 62], feature/channel-level similarity [26, 24, 53], and class-level similarity [48, 54, 56, 55, 57]. Remarkable works have emerged in instance-level methods. For example, RelationKD [45] distills the distance and anglewise correlations between features of different samples. [25] adopts a similar approach by distilling the correlation of samples within a batch through the inner-product calculation of features of each sample. [46] leverages an instance relationship graph where features are represented as nodes and relations as edges for each sample. CIRKD [59], a popular method, extends instance similarity from the batch-level by utilizing a memory bank to consider global relations between samples. [51] conducts similar work by distilling pairwise similarity relations based on stored features to reveal more complete instance relations.
@@
In feature/channel-level similarity, IFVD [26] and ICKD [24] are two pioneering works. The former defines a prototype for each class and distills the distances of pixels from each prototype, while the latter creates a channel correlation matrix using the dot-product of a feature map with its transpose for channel-level similarity distillation. [53] compels the student to imitate the teacher by minimizing the distance between the channel correlation maps of the student and the teacher.
@@
In class-level similarity, DistKD [55] is a popular method that significantly enhances the student’s performance by distilling inter- and intra-class similarities. [54] distills the pairwise similarity of classes by multiplying the logits with its reshaped version. [56] suggests distilling position information and inter-class distances for semantic segmentation. AICSD [57] distills inter-class similarities by introducing intra-class distributions for each class and subsequently calculating pairwise similarity of these distributions using KL for semantic segmentation.
@@
In addition to the similarity levels mentioned, spatial similarity is also utilized for distilling pixel-to-pixel or pixel-to-region [58, 59, 87, 60, 61, 63] similarities. [58] was among the first to propose pooling features into nine regions and distilling the similarity of these regions. CIRKD [59] focused on pixel-to-pixel and pixel-to-region similarities across entire images. [60] transfers the multi-scale pixel-region relation, [61] distills correlations between adjacent blocks of the teacher, [87] employs patch-level similarity for distillation, and [63] introduces distillation of the logits’ similarity from local and global perspectives.
III Schemes
@@ III-A Offline Distillation
Most existing distillation methods follow this offline approach, and the majority of the studies mentioned in the previous section adopt this technique. Offline distillation was first introduced in [10], and numerous methods have since been proposed to extend and improve it. However, offline distillation has some limitations. One major challenge is the memory and computational overhead required to load a large teacher model, which can be a significant issue in resource-constrained environments. Additionally, since the teacher model remains fixed, the student model’s performance is inherently constrained by the teacher’s capabilities. Any biases present in the teacher can also be transferred to the student, potentially affecting its generalization [29, 55].
@@
To address these issues, incorporating adaptive distillation techniques alongside offline distillation can help improve student performance [88, 67, 57]. Moreover, in LLM distillation, where the teacher model is proprietary and only accessible via an API, offline distillation remains the only practical approach. In summary, despite its challenges, offline distillation continues to be the most widely used method in knowledge distillation due to its simplicity and effectiveness.
III-B Online Distillation
@@
Unlike offline distillation, the training of online distillation occurs in a single stage. [89] showed that online distillation offers advantages by using a group of students as a virtual teacher, eliminating the need for a large pretrained model. Early approaches, such as [90] and [91], utilized an ensemble of multi-branch logits to create an on-the-fly teacher and trained the student accordingly. In addition, [92] extended this approach by using the mean of previous predictions as an auxiliary guide to enhance learning. [93] introduced variation in branch depth, with each Hourglass network mimicking both the final heatmap and the fused heatmap of all branches. [94] proposed a framework for mutual contrastive learning, where contrastive embeddings from the same or different networks were used with the anchor during the training phase.
@@
Recent works have focused on improving other aspects of online KD. [95] explores methods to enhance KD by reducing the sharpness and variance gap in the logit space between the teacher and student. [96] aims to achieve flatter loss minima to improve the generalization and stability of student models. [97] enhances multi-task networks by leveraging single-task networks as teachers and employing adaptive feature distillation with online task weighting. Similarly, [98] applies this approach to multi-domain tasks using multiple teachers, where each teacher model specializes in a specific domain.
III-C Self-distillation
@@
Self-distillation is a unique form of online distillation where there is no larger pre-trained teacher; instead, the teacher and student are identical networks [99, 100, 101, 102, 103, 104, 105]. The concept was initially introduced in [99], where a teacher is initially trained on labels, and subsequently, a new identical model is initialized from a different random seed and trained under the guidance of the former teacher. In a similar approach, [101] first trains the network with labels and then utilizes labels and features of the model trained so far for subsequent training steps. [105] incorporates a linear combination of hard targets and predictions from the model in the previous epoch (teacher) to train the network in the current epoch (student). [100] partitions the model into several sections and transfers knowledge from deeper sections to shallower ones, while [104] employs multiple shallow classifiers at different layers and transfers knowledge from the deepest classifier to shallower ones. [103] is the first to provide a theoretical analysis of self-distillation, and [106] establishes a theoretical link between self-distillation and label smoothing.
@@
Recently, a variety of diverse self-distillation algorithms have been proposed [107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117], each offering a new perspective on the problem. Some methods leverage the deeper layers of the model as teachers to distill knowledge to shallower parts of the model [116, 117]. Other approaches utilize predictions from previous epochs as teachers to supervise the model in subsequent epochs [115, 114, 113]. [112] and [110] employ iterative pruning and discarding of parts of the model to create a student for self-distillation. [111] and [108] integrate the concept of self-distillation with graphs, while [109] combines data augmentation, deep contrastive learning, and self-distillation. In this method, different views of the same sample are input to the model for enhanced learning.
IV Algorithms
@@
IV-A Attention-based Distillation
@@
The first work exploring attention distillation was Attention Transfer (AT) [23], which proposed the aggregation of feature channels in the student and teacher models. By minimizing the loss between attention maps of the teacher and student, the attention of the teacher can be effectively transferred to the student. [54] transfers residual attention maps of features from two different layers, while [118] employs a self-attention module to adaptively aggregate context information from the student and teacher feature maps. [34] uses self-attention to simplify the learning process by softening the logits before distillation.
@@
Several methods have focused on utilizing Class Activation Maps (CAM) [119] or Gradient-Weighted Class Activation Maps (GradCAM) [120] for attention distillation by transferring regions with more attention [121, 122, 123]. [124] applies channel self-attention and position self-attention on the features for distillation, [125] employs different types of attention to transfer knowledge at various levels of deep representation learning for KD, and [126] computes attention using student features as queries and teacher features as key values, implementing sparse attention values through random deactivation. These sparse attention values are then used to reweight the feature distance of each teacher-student feature pair to prevent negative transfer.
@@
Recently, powerful attention distillation methods have been proposed [127, 72, 74]. DiffKD [72] suggests training a diffusion model on the teacher’s features and using it to denoise the student’s features to highlight important areas for distillation. AttnFD [74] investigates the role of attention in distillation for semantic segmentation and employs CBAM [75] attention mechanism to refine features before distillation by considering both channel and spatial attention. This refinement focuses on transferring crucial information and emphasizes important regions, particularly enhancing the student network’s performance in scenarios involving small objects.
IV-B Adversarial Distillation
@@
A promising direction for leveraging teacher information during distillation is to embed the process within an adversarial framework inspired by the min-max game of Generative Adversarial Networks (GANs) [128]. Three distinct approaches emerge in this context. First, generative networks can be employed to synthesize the teacher’s training data when the original data is unavailable. Second, a discriminator is integrated to provide refined guidance by distinguishing between teacher and student outputs during knowledge transfer. Third, adversarial training is applied to minimize discrepancies between teacher and student models when confronted with perturbed samples, thus pushing the student to learn from more challenging examples. Collectively, these methods aim to enhance the performance and robustness of the distilled student model, and the following sections will elaborate on these three approaches in detail.
IV-B1 Adversarial Data-Free Knowledge Distillation
@@
Adversarial distillation is applied in various ways, depending on the goal of the distillation process. One approach involves data-free distillation [129, 130, 131, 132, 133, 134, 135, 136, 137, 138, 139, 140, 141], where adversarial examples are generated in the absence of access to the teacher’s original training data, often due to privacy concerns, transmission issues, or legal constraints [131]. Adversarial samples are synthesized either to mimic the original dataset or to augment it, thereby enhancing the distillation process. Methods in this category [129, 132, 135, 130, 137, 140] often leverage the pre-trained teacher network and its internal statistics as a discriminator to guide the generator. Other techniques address distribution shifts in generated samples by proposing the use of an Exponential Moving Average (EMA) of the student as a more reliable learning source [136] or framing the process as a meta-learning problem that balances knowledge acquisition and retention [134]. Self-adversarial strategies [138] have also been explored to extend robustness in data-free settings. Lastly, [133] investigated the challenges of data-free distillation with imbalanced datasets, where the teacher model is pre-trained on skewed class distributions, introducing class-aware strategies to mitigate this imbalance.
IV-B2 Discriminator in the Loop
@@
The use of a discriminator has proven effective in better aligning the prediction distributions between the teacher and student networks. Leveraging this intuition, one of the prevalent approaches in adversarial distillation involves incorporating a discriminator network to align the teacher and student logits [142, 143, 144, 145] or intermediate feature representations [146, 147, 148, 26, 149, 150, 58, 151, 152, 153]. Additionally, some methods employ a three-player framework for discrimination [154, 155], enhancing the effectiveness of the distillation process.
@@
Adversarial distillation has also been extended beyond standard networks to the compression of GANs [156, 157, 158, 159], facilitating the distillation of heavy generator networks into more efficient ones. Notably, the discrimination module is not limited to the GAN framework and has been adapted for diffusion models to reduce the number of inference steps, significantly improving their inference efficiency [160, 161, 162, 163]. Furthermore, advanced techniques have been introduced to utilize the score function instead of the final outputs or logits for distillation, leading to approaches such as adversarial score distillation [164] and variational score distillation [165]. These methods employ score matching and variational approximations to achieve improved robustness and efficiency in knowledge transfer.
IV-B3 Adversarial Robust Knowledge Distillation

### v2

@@ 2.3 Similarity-based Distillation
One of the pioneering works was Yim et al. (2017), which introduced the Flow of Solution Procedure (FSP). This method involved computing the inner product between features from two layers in the teacher network and defining the squared L2 norm as the cost function for each pair of layers between the teacher and student networks. In a related work, You et al. (2017) suggested distilling the relative dissimilarity between intermediate representations of different examples. Subsequent studies proposed distilling the 8-neighborhood similarity of logits (Xie et al., 2018) and distilling similarity of features and logits among multiple teachers (Zhang & Peng, 2018).
@@
Recently, similarity distillation has garnered increased attention, with numerous methods proposed, each focusing on different levels of similarity such as instance/sample-level similarity (Park et al., 2019; Tung & Mori, 2019; Liu et al., 2019f; Peng et al., 2019; Chen et al., 2020b; Cheng et al., 2024a; Hu et al., 2024b; Wang et al., 2024g; Xin et al., 2024), feature/channel-level similarity (Wang et al., 2020d; Liu et al., 2021c; Wang et al., 2023d), and class-level similarity (Chen et al., 2020f; Feng et al., 2021; Zhang et al., 2022d; Huang et al., 2022a; Mansourian et al., 2025; Huang et al., 2025a; Wang et al., 2025a). Remarkable works have emerged in instance-level methods. For example, RelationKD (Park et al., 2019) distills the distance and anglewise correlations between features of different samples. Tung & Mori (2019) adopts a similar approach by distilling the correlation of samples within a batch through the inner-product calculation of features of each sample. Liu et al. (2019f) leverages an instance relationship graph where features are represented as nodes and relations as edges for each sample. CIRKD (Yang et al., 2022c), a popular method, extends instance similarity from the batch-level by utilizing a memory bank to consider global relations between samples. Hu et al. (2024b) conducts similar work by distilling pairwise similarity relations based on stored features to reveal more complete instance relations.
@@
In feature/channel-level similarity, IFVD (Wang et al., 2020d) and ICKD (Liu et al., 2021c) are two pioneering works. The former defines a prototype for each class and distills the distances of pixels from each prototype, while the latter creates a channel correlation matrix using the dot-product of a feature map with its transpose for channel-level similarity distillation. Wang et al. (2023d) compels the student to imitate the teacher by minimizing the distance between the channel correlation maps of the student and the teacher.
@@
In class-level similarity, DistKD (Huang et al., 2022a; 2025a) is a popular method that significantly enhances the student’s performance by distilling inter- and intra-class similarities. Feng et al. (2021) distills the pairwise similarity of classes by multiplying the logits with its reshaped version. Zhang et al. (2022d) suggests distilling position information and inter-class distances for semantic segmentation. AICSD (Mansourian et al., 2025) distills inter-class similarities by introducing intra-class distributions for each class and subsequently calculating pairwise similarity of these distributions using KL for semantic segmentation. IKD (Wang et al., 2025a) facilitates knowledge sharing within the same class to ensure consistent predictions.
@@
In addition to the similarity levels mentioned, spatial similarity is also utilized for distilling pixel-to-pixel or pixel-to-region (Liu et al., 2019d; Yang et al., 2022c; Wang et al., 2023a; b; k; Zhang et al., 2024g) similarities. Liu et al. (2019d) was among the first to propose pooling features into nine regions and distilling the similarity of these regions. CIRKD (Yang et al., 2022c) focused on pixel-to-pixel and pixel-to-region similarities across entire images. Wang et al. (2023b) transfers the multi-scale pixel-region relation, Wang et al. (2023k) distills correlations between adjacent blocks of the teacher, Wang et al. (2023a) employs patch-level similarity for distillation, and Zhang et al. (2024g) introduces distillation of the logits’ similarity from local and global perspectives.
3 Schemes
@@ 3.1 Offline Distillation
Most existing distillation methods follow this offline approach, and the majority of the studies mentioned in the previous section adopt this technique. Offline distillation was first introduced in Hinton (2015), and numerous methods have since been proposed to extend and improve it. However, offline distillation has some limitations. One major challenge is the memory and computational overhead required to load a large teacher model, which can be a significant issue in resource-constrained environments. Additionally, since the teacher model remains fixed, the student model’s performance is inherently constrained by the teacher’s capabilities. Any biases present in the teacher can also be transferred to the student, potentially affecting its generalization (Guo et al., 2020a; Huang et al., 2022a).
@@
To address these issues, incorporating adaptive distillation techniques alongside offline distillation can help improve student performance (Cho & Hariharan, 2019; Shu et al., 2021; Mansourian et al., 2025). Moreover, in LLM distillation, where the teacher model is proprietary and only accessible via an API, offline distillation remains the only practical approach. In summary, despite its challenges, offline distillation continues to be the most widely used method in knowledge distillation due to its simplicity and effectiveness.
3.2 Online Distillation
@@
Unlike offline distillation, the training of online distillation occurs in a single stage (Chen et al., 2020a). Wang et al. (2024a) showed that online distillation offers advantages by using a group of students as a virtual teacher, eliminating the need for a large pretrained model. Early approaches, such as Guo et al. (2020b) and Zhu et al. (2018), utilized an ensemble of multi-branch logits to create an on-the-fly teacher and trained the student accordingly. In addition, Wu & Gong (2021) extended this approach by using the mean of previous predictions as an auxiliary guide to enhance learning. Li et al. (2021b) introduced variation in branch depth, with each Hourglass network mimicking both the final heatmap and the fused heatmap of all branches. Yang et al. (2023b) proposed a framework for mutual contrastive learning, where contrastive embeddings from the same or different networks were used with the anchor during the training phase.
@@
Recent works have focused on improving other aspects of online KD. Rao et al. (2023) explores methods to enhance KD by reducing the sharpness and variance gap in the logit space between the teacher and student. Zhang et al. (2023h) aims to achieve flatter loss minima to improve the generalization and stability of student models. Jacob et al. (2023) enhances multi-task networks by leveraging single-task networks as teachers and employing adaptive feature distillation with online task weighting. Similarly, Ypsilantis et al. (2024) applies this approach to multi-domain tasks using multiple teachers, where each teacher model specializes in a specific domain.
3.3 Self-distillation
@@
Self-distillation is a unique form of online distillation where there is no larger pre-trained teacher; instead, the teacher and student are identical networks (Furlanello et al., 2018; Zhang et al., 2019b; Lan et al., 2019; Xu & Liu, 2019; Mobahi et al., 2020; Zhang et al., 2021c; Kim et al., 2021a). The concept was initially introduced in Furlanello et al. (2018), where a teacher is initially trained on labels, and subsequently, a new identical model is initialized from a different random seed and trained under the guidance of the former teacher. In a similar approach, Lan et al. (2019) first trains the network with labels and then utilizes labels and features of the model trained so far for subsequent training steps. Kim et al. (2021a) incorporates a linear combination of hard targets and predictions from the model in the previous epoch (teacher) to train the network in the current epoch (student). Zhang et al. (2019b) partitions the model into several sections and transfers knowledge from deeper sections to shallower ones, while Zhang et al. (2021c) employs multiple shallow classifiers at different layers and transfers knowledge from the deepest classifier to shallower ones. Mobahi et al. (2020) is the first to provide a theoretical analysis of self-distillation, and Zhang & Sabuncu (2020) establishes a theoretical link between self-distillation and label smoothing.
@@
Recently, a variety of diverse self-distillation algorithms have been proposed (Yang et al., 2023g; Zhang et al., 2023b; Xiao et al., 2023b; Zheng et al., 2024b; Wu et al., 2024a; Chen et al., 2024a; Lin et al., 2024b; Li et al., 2024i; Liu et al., 2024e; Liang et al., 2024; Qin et al., 2025; Wang et al., 2025c), each offering a new perspective on the problem. Some methods leverage the deeper layers of the model as teachers to distill knowledge to shallower parts of the model (Liang et al., 2024; Qin et al., 2025). Other approaches utilize predictions from previous epochs as teachers to supervise the model in subsequent epochs (Liu et al., 2024e; Li et al., 2024i; Lin et al., 2024b). Chen et al. (2024a) and Zheng et al. (2024b) employ iterative pruning and discarding of parts of the model to create a student for self-distillation and MSKD(Wang et al., 2025c) proposes a self-distillation method for multi-label learning. Wu et al. (2024a) and Zhang et al. (2023b) integrate the concept of self-distillation with graphs, while Xiao et al. (2023b) combines data augmentation, deep contrastive learning, and self-distillation. In this method, different views of the same sample are input to the model for enhanced learning.
4 Algorithms
@@
4.1 Attention-based Distillation
@@
The first work exploring attention distillation was Attention Transfer (AT) (Zagoruyko & Komodakis, 2016), which proposed the aggregation of feature channels in the student and teacher models. By minimizing the loss between attention maps of the teacher and student, the attention of the teacher can be effectively transferred to the student. Feng et al. (2021) transfers residual attention maps of features from two different layers, while An et al. (2022) employs a self-attention module to adaptively aggregate context information from the student and teacher feature maps. Yuan et al. (2024b) uses self-attention to simplify the learning process by softening the logits before distillation.
@@
Several methods have focused on utilizing Class Activation Maps (CAM) (Zhou et al., 2016) or Gradient-Weighted Class Activation Maps (GradCAM) (Selvaraju et al., 2017) for attention distillation by transferring regions with more attention (Bavandpour & Kasaei, 2020; Cho & Kang, 2022; Guo et al., 2023d). Karine et al. (2024) applies channel self-attention and position self-attention on the features for distillation, Gou et al. (2023) employs different types of attention to transfer knowledge at various levels of deep representation learning for KD, and Guo et al. (2024b) computes attention using student features as queries and teacher features as key values, implementing sparse attention values through random deactivation. These sparse attention values are then used to reweight the feature distance of each teacher-student feature pair to prevent negative transfer.
@@
Recently, powerful attention distillation methods have been proposed (Yang et al., 2023d; Huang et al., 2023b; Mansourian et al., 2024). DiffKD (Huang et al., 2023b) suggests training a diffusion model on the teacher’s features and using it to denoise the student’s features to highlight important areas for distillation. AttnFD (Mansourian et al., 2024) investigates the role of attention in distillation for semantic segmentation and employs CBAM (Woo et al., 2018) attention mechanism to refine features before distillation by considering both channel and spatial attention. This refinement focuses on transferring crucial information and emphasizes important regions, particularly enhancing the student network’s performance in scenarios involving small objects.
4.2 Adversarial Distillation
@@
A promising direction for leveraging teacher information during distillation is to embed the process within an adversarial framework inspired by the min-max game of Generative Adversarial Networks (GANs) (Goodfellow et al., 2014). Three distinct approaches emerge in this context. First, generative networks can be employed to synthesize the teacher’s training data when the original data is unavailable. Second, a discriminator is integrated to provide refined guidance by distinguishing between teacher and student outputs during knowledge transfer. Third, adversarial training is applied to minimize discrepancies between teacher and student models when confronted with perturbed samples, thus pushing the student to learn from more challenging examples. Collectively, these methods aim to enhance the performance and robustness of the distilled student model, and the following sections will elaborate on these three approaches in detail.
4.2.1 Adversarial Data-Free Knowledge Distillation
@@
Adversarial distillation is applied in various ways, depending on the goal of the distillation process. One approach involves data-free distillation (Chen et al., 2019a; Ye et al., 2020; Fang et al., 2019; Choi et al., 2020; Liao et al., 2024; Patel et al., 2023; Wang et al., 2024o; Do et al., 2022; Zhao et al., 2022b; Zhou et al., 2024d; Liu et al., 2018; Micaelli & Storkey, 2019; Zhang et al., 2022b), where adversarial examples are generated in the absence of access to the teacher’s original training data, often due to privacy concerns, transmission issues, or legal constraints (Fang et al., 2019). Adversarial samples are synthesized either to mimic the original dataset or to augment it, thereby enhancing the distillation process. Methods in this category (Chen et al., 2019a; Choi et al., 2020; Wang et al., 2024o; Ye et al., 2020; Zhao et al., 2022b; Micaelli & Storkey, 2019) often leverage the pre-trained teacher network and its internal statistics as a discriminator to guide the generator. Other techniques address distribution shifts in generated samples by proposing the use of an Exponential Moving Average (EMA) of the student as a more reliable learning source (Do et al., 2022) or framing the process as a meta-learning problem that balances knowledge acquisition and retention (Patel et al., 2023). Self-adversarial strategies (Zhou et al., 2024d) have also been explored to extend robustness in data-free settings. Lastly, Liao et al. (2024) investigated the challenges of data-free distillation with imbalanced datasets, where the teacher model is pre-trained on skewed class distributions, introducing class-aware strategies to mitigate this imbalance.
4.2.2 Discriminator in the Loop
@@
The use of a discriminator has proven effective in better aligning the prediction distributions between the teacher and student networks. Leveraging this intuition, one of the prevalent approaches in adversarial distillation involves incorporating a discriminator network to align the teacher and student logits (Xu et al., 2017; Liu et al., 2019a; Ghorbani et al., 2020; Croitoru et al., 2024) or intermediate feature representations (Liu et al., 2020a; Wang et al., 2020a; Chung et al., 2020; Wang et al., 2020d; Zhang et al., 2020b; Li et al., 2020b; Liu et al., 2019d; Shen et al., 2019d; Belagiannis et al., 2018; Wang et al., 2018c). Additionally, some methods employ a three-player framework for discrimination (Wang et al., 2019; 2018b), enhancing the effectiveness of the distillation process.
@@
Adversarial distillation has also been extended beyond standard networks to the compression of GANs (Aguinaldo et al., 2019; Wang et al., 2020c; Li et al., 2020a; Zhang et al., 2022a), facilitating the distillation of heavy generator networks into more efficient ones. Notably, the discrimination module is not limited to the GAN framework and has been adapted for diffusion models to reduce the number of inference steps, significantly improving their inference efficiency (Sauer et al., 2024; Liu et al., 2024a; Kong et al., 2024; Lin et al., 2024c). Furthermore, advanced techniques have been introduced to utilize the score function instead of the final outputs or logits for distillation, leading to approaches such as adversarial score distillation (Wei et al., 2024a) and variational score distillation (Wan et al., 2024). These methods employ score matching and variational approximations to achieve improved robustness and efficiency in knowledge transfer.
4.2.3 Adversarial Robust Knowledge Distillation